# CloudTrack: Scalable UAV Tracking with Cloud Semantics

### v2

@@ Abstract
Nowadays, unmanned aerial vehicles (UAVs) are commonly used in search and rescue scenarios to gather information in the search area. The automatic identification of the person searched for in aerial footage could increase the autonomy of such systems, reduce the search time, and thus increase the missed person’s chances of survival. In this paper, we present a novel approach to perform semantically conditioned open vocabulary object tracking that is specifically designed to cope with the limitations of UAV hardware. Our approach has several advantages. It can run with verbal descriptions of the missing person, e.g., the color of the shirt, it does not require dedicated training to execute the mission and can efficiently track a potentially moving person. Our experimental results demonstrate the versatility and efficacy of our approach. We publish the source code of our approach at https://github.com/utn-blei/CloudTrack.
@@ I Introduction
Visual detection and tracking of objects is an integral part of many modern-day UAV applications. Traditional object detection approaches [1, 2, 3] based on Convolutional Neural Networks (CNNs) require careful collection and curation of data to reach acceptable levels of performance. In search and rescue (SAR) operations, the object of interest is usually unknown when developing the system. Thus, traditional approaches are often inapplicable or limited to a small set of pre-trained classes. Instead, in many SAR scenarios, a semantic description is given, e.g., “The lost person is male and wears a gray shirt” or “The person was last seen driving in a red car near a forest”. Even though modern Vision-Language Models (VLMs) can exploit such information, they are typically inapplicable in UAVs due to a limitation of onboard hardware resources. Further, comparably low frame rates of such pipelines limit the ability to track objects of interest across multiple frames.
@@ II-A Detection and Tracking for UAVs
Numerous approaches for person detection and tracking in SAR missions have been proposed. The focus often lies on enabling on-board processing, for instance, by deploying lightweight models [7, 8, 9] like YOLO [1]. Xu et al. [10] tackle the problem of limited edge computing with a collaborative learning scheme. A lightweight onboard network sends noisy person detections to the mission center for verification. Using the corresponding feedback, the UAV network is improved over time.
A similar idea is used by Zhou and Liu [11] to improve an on-board person tracker through human operator feedback provided by a gaze tracking system. Our method also relies on a connection to a ground station for tracking initialization but we deploy a VLM instead of a human operator. Generally, detection and tracking can be both performed on an edge device. For instance, Lo et al. [12] deploy YOLO for detection and a Kalman filter for general object tracking. Similarly, Chen et al. [13] detect pedestrians with YOLO and propose a motion model-based tracking algorithm. Both methods deliver acceptable performance on closed, pre-trained object sets. However, current drone computing is too restrictive for full open-world detection.
@@ III-A Overview
III-B Object Detection Back-End
The back end consists of an OV object detector and a VLM. Notably, the OV object detector evaluated in this work demonstrates strong performance even when identifying small instances relative to the overall image size. It struggles, however, to understand the deeper semantics of a prompt as it is common in SAR missions. Thus, prompting the system with, e.g., "An injured male person wearing a gray shirt.", results in bounding boxes of all individuals in the image, independent of potential signs of injury, a match in cloth color or gender. This leads to a high number of false positives and could overwhelm human supervisors. VLMs on the other hand feature strong capabilities of semantic understanding. Contrary, they struggle to correctly identify small objects in the image. This leads to unexpected behavior, such as random answers and hallucinations.
@@ III-C Object Tracking Front-End
In SAR missions, targets are often in motion, such as lost children or individuals involved in maritime accidents. Therefore, a tracking system is essential to maintain an accurate, real-time position of any potential detection. A naïve approach is to stream every image from the UAV to the back-end and complete tracking on the back-end platform. In many cases, data links are of limited bandwidth and potentially unstable. This leads to slow and unpredictable tracking frame rates, high delays in UAV course corrections and thus to the potential loss of an object of interest.
We, therefore, require an on-board solution for short-term object tracking. This solution must support the OV nature of our pipeline. We further require the system to run with an adequate frame rate of at least on common drone companion computers from the Nvidia Jetson and Raspberry Pi series. To this end, we benchmark six compatible lightweight trackers on common drone hardware (Raspberry Pi 4 (2 Gb) and NVIDIA Jetson Orin NX (16 Gb)). We evaluate each tracker on our two platforms and then select the best-performing tracker in terms of tracking accuracy.
@@ IV-A Search and Rescue Image Dataset for Person Detection (SARD)
The “Search and Rescue Image Dataset for Person Detection (SARD)” [6] consists of 6,532 instances of actors in different poses on 1,981 images. The images are taken from a UAV perspective in a natural setting. The dataset natively contains ground truth for the actors bounding boxes and poses (i.e., standing, lying, etc.). As part of this work, we annotate the color of each person’s shirt. We approach this in a two-step manner. First, we prompt the GPT-4o API with a cropped image patch, containing the individual, for the color of the shirt. In the second step, we manually review the results for each individual. We then add ground truth for the recommendation for help. We first assume that all individuals requiring assistance are instances of the pose classes "laying_down", "not_defined", "null" and "seated". In many cases, classification based on the pose alone is not sufficient. We, therefore, classify all resulting 3,255 instances individually. We label individuals as “injured”, if they are in a pose signalizing pain, injury, or unconsciousness.
@@ V-B OV Detection Evaluation
First, we design four types of detection tasks suitable for SAR missions, each constituting a person’s description with a different level of semantic complexity. By varying the person’s cloth color and posture, we in total obtain 8 referring expressions as detection objectives:
@@
We compare our approach with GPT4-mini as VLM against GLEE [28] and Grounding DINO, two state-of-the-art OV object detection methods. For GLEE, we evaluate the versions with the largest backbone, GLEE plus and GLEE pro. Our method also uses Grounding DINO for initial object detection but with an additional verification step by a VLM. We evaluate the baselines on the same data and tasks as above. As displayed in table II, none of the tested baselines can achieve a similar semantic understanding accuracy as our approach. Grounding DINO as our best-tested baseline performs comparably well on the task of detecting a person. It struggles, however, when prompted for semantically more complex scenarios. Further, its performance is not stable across several evaluations of the same semantic complexity. This becomes clear when comparing mAP scores on the “shirt_gray” experiment with the performance on the “shirt_blue” experiment. According to the authors, GLEE shows strong performance when tracking an object across several frames. But in their evaluation, the authors prompt the system with less semantically complex scenarios than used in our comparison. We believe this is a possible reason for its sub-average performance in our experiments.
@@ V-C Full-Pipeline Evaluation
We next analyze the performance of our complete pipeline on two UAV companion computers and hardware set-up as described in Section V-A. As the back-end, we deploy the configuration using LLaVA13b as VLM. Our main goal is to demonstrate that our OV tracking approach can run in real-time on this limited hardware. In the process, we also benchmark several tracking front-end choices on both companion platforms. We conduct the evaluation on the Referring VOT22 benchmark (see Section IV-B) to obtain use case agnostic numbers due to a wider range of object classes and more complex object trajectories compared to SARD.
@@ VI Conclusion and Future Work
In this paper, we introduce CloudTrack, a scalable OV object tracking approach with advanced semantic understanding. Our approach combines powerful foundation models in a cloud back-end with a real-time capable tracker on the UAVs companion computer. We extensively evaluate our framework on two common UAV companion computers, four different VLMs, and six object trackers to find the best-suited configuration for each hardware platform. As we show in our evaluation, our framework significantly outperforms other approaches in terms of semantic understanding. This, however, comes at the cost of increased runtime in the back end. Future work could tackle this problem by fine-tuning the OV object detector or combining the results of fast and slow VLMs to reduce the latency time of the UAV. Further, multi-object tracking could be supported to permit use cases beyond SAR missions with more than one object of interest.

### v3

@@ Abstract
Nowadays, unmanned aerial vehicles (UAVs) are commonly used in search and rescue scenarios to gather information in the search area. The automatic identification of the person searched for in aerial footage could increase the autonomy of such systems, reduce the search time, and thus increase the missed person’s chances of survival. In this paper, we present a novel approach to perform semantically conditioned open vocabulary object tracking that is specifically designed to cope with the limitations of UAV hardware. Our approach has several advantages: It can run with verbal descriptions of the missing person, e.g., the color of the shirt, it does not require dedicated training to execute the mission, and can efficiently track a potentially moving person. Our experimental results demonstrate the versatility and efficacy of our approach. We publish the methods source code at https://github.com/yblei/CloudTrack.
@@ I INTRODUCTION
Visual detection and tracking of objects is an integral part of many modern-day UAV applications. Traditional object detection approaches [1, 2, 3] based on Convolutional Neural Networks (CNNs) require careful collection and curation of data to reach acceptable levels of performance. In search and rescue (SAR) operations, the object of interest is usually unknown when developing the system. Thus, traditional approaches are often inapplicable or limited to a small set of pre-trained classes. Instead, in many SAR scenarios, a semantic description is given, e.g., “The lost person is male and wears a gray shirt” or “The person was last seen driving in a red car near a forest”. Even though modern Vision Language Models (VLMs) can exploit such information, they are typically inapplicable in UAVs due to a limitation of onboard hardware resources. Further, comparably low frame rates of such pipelines limit the ability to track objects of interest across multiple frames.
@@ II-A Detection and Tracking for UAVs
Numerous approaches for person detection and tracking in SAR missions have been proposed. The focus often lies on enabling onboard processing, for instance, by deploying lightweight models [7, 8, 9] like YOLO [1]. Xu et al. [10] tackle the problem of limited edge computing with a collaborative learning scheme. A lightweight onboard network sends noisy person detections to the mission center for verification. Using the corresponding feedback, the UAV network is improved over time.
A similar idea is used by Zhou and Liu [11] to improve an onboard person tracker through human operator feedback provided by a gaze tracking system. Our method also relies on a connection to a ground station for tracking initialization but we deploy a VLM instead of a human operator. Generally, detection and tracking can be both performed on an edge device. For instance, Lo et al. [12] deploy YOLO for detection and a Kalman filter for general object tracking. Similarly, Chen et al. [13] detect pedestrians with YOLO and propose a motion model-based tracking algorithm. Both methods deliver acceptable performance on closed, pre-trained object sets. However, current drone computing is too restrictive for full open-world detection.
@@ III-A Overview
III-B Object-Detection Back-End
The back-end consists of an OV object detector and a VLM. Notably, the OV object detector evaluated in this work demonstrates strong performance even when identifying small instances relative to the overall image size. It struggles, however, to understand the deeper semantics of a prompt as it is common in SAR missions. Thus, prompting the system with, e.g., "An injured male person wearing a gray shirt.", results in bounding boxes of all individuals in the image, independent of potential signs of injury, a match in cloth color or gender. This leads to a high number of false positives and could overwhelm human supervisors. VLMs on the other hand feature strong capabilities of semantic understanding. On the contrary, they struggle to correctly identify small objects in the image. This leads to unexpected behavior, such as random answers and hallucinations.
@@ III-C Object Tracking Front-End
In SAR missions, targets are often in motion, such as lost children or individuals involved in maritime accidents. Therefore, a tracking system is essential to maintain an accurate, real-time position of any potential detection. A naïve approach is to stream every image from the UAV to the back-end and complete tracking on the back-end platform. Given adequate resources, this could either be the ground control station or a dedicated cloud server. In many cases, data links are of limited bandwidth and potentially unstable. This leads to slow and unpredictable tracking frame rates, high delays in UAV course corrections and thus to the potential loss of an object of interest.
We, therefore, require an onboard solution for short-term object tracking. This solution must support the OV nature of our pipeline. We further require the system to run with an adequate frame rate of at least on common drone companion computers from the NVIDIA Jetson and Raspberry Pi series. To this end, we benchmark six compatible lightweight trackers on common drone hardware (Raspberry Pi 4 (2 Gb) and NVIDIA Jetson Orin NX (16 Gb)). We evaluate each tracker on our two platforms and then select the best-performing tracker in terms of tracking accuracy.
@@ IV-A Search and Rescue Image Dataset for Person Detection (SARD)
The “Search and Rescue Image Dataset for Person Detection (SARD)” [6] consists of instances of actors in different poses on images. The images are taken from a UAV perspective in a natural setting. The dataset natively contains ground truth for the actors bounding boxes and poses (i.e., standing, lying, etc.). As part of this work, we annotate the color of each person’s shirt. We approach this in a two-step manner. First, we prompt the GPT-4o API with a cropped image patch, containing the individual, for the color of the shirt. In the second step, we manually review the results for each individual. We then add ground truth for the recommendation for help. We first assume that all individuals requiring assistance are instances of the pose classes "laying_down", "not_defined", "null" and "seated". In many cases, classification based on the pose alone is not sufficient. We, therefore, classify all resulting instances individually. We label individuals as "injured", if they are in a pose signalizing pain, injury, or unconsciousness.
@@ V-B OV Detection Evaluation
First, we design four types of detection tasks suitable for SAR missions, each constituting a person’s description with a different level of semantic complexity. By varying the person’s cloth color and posture, we in total obtain referring expressions as detection objectives:
@@
We compare our approach with GPT4-mini as VLM against GLEE [28] and Grounding DINO, two state-of-the-art OV object detection methods. For GLEE, we evaluate the versions with the largest backbone, GLEE plus and GLEE pro. Our method also uses Grounding DINO for initial object detection but with an additional verification step by a VLM. We evaluate the baselines on the same data and tasks as above. As displayed in Table II, none of the tested baselines can achieve a similar semantic understanding accuracy as our approach. Grounding DINO as our best-tested baseline performs comparably well on the task of detecting a person. It struggles, however, when prompted for semantically more complex scenarios. Further, its performance is not stable across several evaluations of the same semantic complexity. This becomes clear when comparing mAP scores on the "shirt_gray" experiment with the performance on the "shirt_blue" experiment. According to the authors, GLEE shows strong performance when tracking an object across several frames. But in their evaluation, the authors prompt the system with less semantically complex scenarios than used in our comparison. We believe this is a possible reason for its sub-average performance in our experiments.
@@ V-C Full-Pipeline Evaluation
We next analyze the performance of our complete pipeline on two UAV companion computers and hardware set-up as described in Section V-A. As the back-end, we deploy the configuration using LLaVA13b as VLM. Our main goal is to demonstrate that our OV tracking approach can run in real-time on this limited hardware. In the process, we also test several tracking front-end choices on both companion platforms. We conduct the evaluation on the Referring VOT22 benchmark (see Section IV-B) to obtain use case agnostic numbers due to a wider range of object classes and more complex object trajectories compared to SARD.
@@ VI CONCLUSION AND FUTURE WORK
In this paper, we introduce CloudTrack, a scalable OV object tracking approach with advanced semantic understanding. Our approach combines powerful foundation models in a cloud back-end with a real-time capable tracker on the UAVs companion computer. We extensively evaluate our framework on two common UAV companion computers, four different VLMs, and six object trackers to find the best-suited configuration for each hardware platform. As we show in our evaluation, our framework significantly outperforms other approaches in terms of semantic understanding. This, however, comes at the cost of increased runtime in the back-end. Future work could tackle this problem by fine-tuning the OV object detector or combining the results of fast and slow VLMs to reduce the latency time of the UAV. Further, multi-object tracking could be supported to permit use cases beyond SAR missions with more than one object of interest.